Mattress of a water bed is 2.0m long by 2.0m wide and 0.3m.
Find the weight of water in the mattress(density of water is 100kg/m3.
Find the pressure exerted by water bed on the flow when the bed rests on it's normal position.assume the entire longer surface of bed makes contact with the flow.
1
Expert's answer
2021-08-03T11:41:04-0400

The volume of the mattress

"V=\rm 2.0\: m\times 2.0\: m\times 0.3\: m=1.2\: m^3"

The mass of water in the mattress

"m=\rho V=\rm 1000\: rg/m^3\times1.2\: m^3=1200\: kg"

The pressure

"P=\frac{mg}{A}=\frac{1200*9.8}{2.0*2.0}=2940\:\rm Pa"